An RL series circuit is constructed with a 133 Ω resistor and a 58 mH inductor. A generator with a frequency of 632 Hz and an rms voltage of 45 V provides the voltage for the circuit. What is the predicted phase angle in degrees for the total rms voltage? Input the answer in scientific notation with “E” notation. For example, 39.85 rounded to two significant figures would be entered as 4.0E1.
Answer #1

Step 1: Find inductive reactance of given inductor:

XL = w*L = 2*pi*f*L

f = frequency of circuit = 632 Hz

L = Inductance = 58 mH

So, XL = 2*pi*632*58*10^-3 = 230.32 ohm

R = Resistance = 133 ohm

Step 2: Now phase angle in RL circuit is given by:

= arctan (XL/R)

= arctan (230.32/133) = 59.995 deg

In two significant figures

phase angle = 60 deg = 6.0E1 degree

Let me know if you've any query.